OpenStudy (anonymous):

Can anyone help me with this problem. Find an equation of the given pair of point.s Express answer in form of x=a,y=b, or y=mx+b (-7,-7) and (5,5)

OpenStudy (anonymous):

First you need to find the slope

OpenStudy (anonymous):

(5 - -7)/(5 - -7) (5 + 7)(5 + 7) 12/12 1 So slope = 1

OpenStudy (anonymous):

so the equation is f(x)=x

OpenStudy (anonymous):

Then you use this formula (y - y.) = slope(x - x.) My "." usually have a subscript of 1 but I don't know how to make that happen. The y. and the x. are the x's and y's that you you know. You pick one point and use those values. I am going to use the (5,5) point

OpenStudy (anonymous):

(y - 5) = 1(x - 5) y - 5 = x - 5 add 5 to both sides y = x This is your answer